# NIMO: a Nonlinear Interpretable MOdel

### v1

@@ 3.2 Training through profile likelihood with adaptive ridge regression
Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\bm{y}\in\mathbb{R}^{d}$. As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$. The model in Eq. (2) can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication. Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\bm{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$. To disentangle the optimization of $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood (Venzon and Moolgavkar, 1988) and consider an optimization based on $\bm{u}$ only. Since we know the analytical solution of $\bm{\beta}$ for ridge regression, we can alternate one optimization step over $\bm{u}$ with an analytical update over $\bm{\beta}$:
$\min_{\bm{\beta},\bm{u}}\mathcal{L}(\bm{\beta},\bm{u})\quad\rightarrow\quad\min_{\bm{u}}\mathcal{L}(\hat{\bm{\beta}},\bm{u})\quad\text{where}\quad\hat{\bm{\beta}}=(B_{\bm{u}}^{T}B_{\bm{u}}+\lambda I)^{-1}B_{\bm{u}}^{T}\mathbf{y}.$ (3)
In practice, $\hat{\bm{\beta}}$ depends on the current estimate of $\bm{u}$ via $B_{\bm{u}}$ so we alternate one gradient descent step over $\bm{u}$ with one analytical update over $\bm{\beta}$ according to $\hat{\bm{\beta}}$ in Eq. (3). Note that in practice this is much more effective than freely learning via gradient descent $\bm{\beta}$ alongside with $\bm{u}$. Furthermore, this formalism allows to include sparsity by simply replacing the L2 penalty with the L1 penalty:
@@
In contrast to the ridge objective, there is no closed-form solution for Lasso regression. Instead, we can use adaptive ridge regression, which is equivalent to Lasso (Grandvalet, 1998); see the Appendix for a detailed proof. By rewriting our objective as in adaptive ridge regression, we can enforce sparsity and find an analytical update for $\bm{\beta}$. The exact algorithm is shown in Alg. 1 in the Appendix.
@@
The same idea can be extended to any generalized linear model. For simplicity, consider logistic regression with L1 penalty. As before, we first get a closed-form expression for $\bm{\beta}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$. What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate, which is a weighted least-squares approximation of the original problem:
@@
where $W_{ii}=\sigma(\eta_{i})\bigl{[}1-\sigma(\eta_{i})\bigr{]}$ and $\bm{\eta}=B_{\bm{u}}\bm{\beta}$. The so-called working response $\bm{z}$ is computed as $\bm{z}=\bm{\eta}+W^{-1}\bigl{(}\mathbf{y}-\sigma(\bm{\eta})\bigr{)}$. The problem in Eq. (5) is also a lasso problem and we can still employ adaptive ridge regression to find a closed form solution for $\bm{\beta}$. Then we substitute $\bm{\beta}$ back and optimize over the neural networks parameters $\bm{u}$. A detailed derivation can be found in the Appendix.
3.3 Sparsity of nonlinear interactions
In the previous paragraphs we showed that linear coefficients $\bm{\beta}$ in NIMO are interpretable as in a linear model. Even though the neural network contribution is not interpretable per se, we can still identify relevant nonlinear interactions by means of sparsity. The idea is to force the first layer of the neural network to be sparse, hence to use only relevant features. To do so, we apply a group penalty on the weight matrix $W\in\mathbb{R}^{p\times d}$ of the first fully-connected (fc) layer, where $d$ is the number of input features and $p$ the hidden dimension of the first fc layer. In order to obtain sparsity at a input feature level, we need to consider neurons connected to such feature and penalize them as a group:
$\ell_{\text{group}}=\lambda_{\text{group}}\sum_{j=1}^{d}\|\mathbf{w}_{j}\|_{2}\;.$ (6)
where $\mathbf{w}_{j}=W[:,j]\in\mathbb{R}^{p}$ is the weight vector acting on the $j$-th feature. With this penalty on the first fc layer, we encourage it to select only certain input features. However, since an MLP computes a composition of functions, the subsequent fc layers can still compensate for the sparsity induced by the regularization in the first layer. Specifically, even if the first layer produces sparse outputs, subsequent layers can reweigh and transform these outputs to restore their influence on the final prediction. To mitigate this compensation effect, we inject noise into the output of the first layer. By doing so, we disrupt the deterministic signal path, making it harder for the subsequent layers to simply exploit the resulting weak, regularized, and noisy signal. As a result, the network is encouraged to focus on strong and robust signals associated with informative input features.This approach works very well in practice and allows to easily achieve sparsity. We provide further details in the Appendix.
3.4 Limitations
The proposed approach is designed to maintain the interpretability of linear models and, at the same time, to learn nonlinear corrections to linear predictions. Therefore, the underlying assumption is that the data can be described by adding nonlinearities over a linear model. Highly nonlinear datasets might not be suitable for the proposed approach. However, in such cases the question itself about interpretability would be ill-posed, at least according to our notion of interpretability. There is however one case that cannot be modeled by Eq. (2). Assume the underlying data generating process contains a self-interacting term of the kind $x_{j}f(x_{j})$, where $f$ is a linear or nonlinear function (e.g. $x\sin(x)$ or $x^{2}$). In order to make the linear coefficients interpretable we assumed $g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$ not to depend on $x_{j}$, which clearly prevents us from learning such types of nonlinearities. However, if such feature is known to play a role in the dataset, simple feature engineering can solve the issue.
@@ 4 Experiments
In this section, we experimentally showcase the proposed approach. Firstly, in Section 4.1 we use a toy example to illustrate how the model works. Secondly, in Section 4.2 we show that across several synthetic datasets the model learns the actual underlying coefficients, the nonlinearities and the correct sparsity in the networks. We also achieve higher performance compared to other hybrid approaches. Lastly, in Section 4.3 we test the model on real datasets and show that it is able to select a sparser set of features than Lasso, while maintaining consistency with Lasso in terms of the signs and magnitudes of the selected coefficients. However, thanks to the nonlinear terms the model achieves significantly higher performance, on par with neural networks but with the benefits of inherent interpretability. In the Appendix we provide further details on each experimental setup.
4.1 Toy example
We first illustrate NIMO on a simple three-dimensional toy example. We generate the data as follows:
where $\epsilon$ is the usual Gaussian noise. This toy example has 3 features $\{x_{`},x_{2},x_{3}\}$, but $x_{3}$ is actually inactive (hence uninformative) since it is multiplied by a zero coefficient. We generate 400 samples and use a 200-100-100 train-validation-test split. As a baseline, we compare with Lasso regression. As shown in Figure 3 (left) the $\bm{\beta}$ coefficients learned with NIMO coincide with the ground truth. Relevantly, NIMO correctly learns $\beta_{3}$ to be zero, i.e. that $x_{3}$ is uninformative. In Figure 3 (middle) we show that the neural network used in NIMO can perfectly recover the underlying nonlinear interactions both for $x_{1}$ and $x_{2}$. Furthermore, if we analyze the weights of the first fc layer, we can see that the neurons connected to the input $x_{3}$ are sparse; see Figure 3 (right). This provides further interpretability in the model because it implies that $x_{3}$ does not contribute to nonlinear interactions.
4.2 Synthetic Experiments
We now showcase that NIMO is able to recover sparse coefficients and learn the underlying nonlinearities across a set of different regression and classification settings, with different number of features and sparsity levels. We compare out method against Lasso regression (Tibshirani, 1996), a vanilla neural network and two hybrid approaches, namely LassoNet (Lemhadri et al., 2021), Contextual Lasso (Thompson et al., 2023). Contextual Lasso requires the features to be separated in explanatory and contextual, which is not directly applicable to our settings since we do not have contextual features. To still compare with Contextual Lasso, we use all features as context. Compared with other hybrid approaches, NIMO achieves better or equivalent performance while still retaining interpretability. In all settings we use 200 training samples and 100 for validation and test. Further details can be found in the Appendix.
@@ 4.2 Synthetic Experiments
Regression We design synthetic regression datasets by explicitly controlling the linear coefficients and the nonlinearities, similarly to Eq. (7). We explore multiple nonlinearities, different dimensionality and also use several uninformative features (i.e. with zero coefficient) to test if we can recover sparse signals. As a sanity check, we first test NIMO on a Vanilla regression dataset with linear coefficients in $\{-5,-4,...,3,4\}$ but without nonlinearities. The idea is to check if the neural network of NIMO interferes with the linear part in the absence of nonlinearities. As shown in Figure 4 (left), NIMO is able to perfectly recover the ground truth coefficients, and the neural network does not interfere with them. We provide three more regression settings (Setting 1, 2, 3) where we test different nonlinearities and different sparsity levels. We provide details about the exact data generating process in the Appendix. In Table 1 we report the MSE on the test set over all settings for NIMO and we compare it against linear, nonlinear and hybrid approaches. Overall, NIMO achieves the best performance across all settings with nonlinearities. This is possible because NIMO learns the correct coefficients (Figure 4) and, at the same time, accurately learns the underlying nonlinear interactions. Differently from Lasso, we can also perfectly learn the sparsity pattern in the coefficients, i.e. which features are uninformative; see Figure 4 (bottom, right). In contrast, nonlinear models and hybrid models learn spurious nonlinear relationships.
Table 1: MSE loss for synthetic regression settings (see the Appendix for more details). We compare with Lasso regression, NN, LassoNet and Contextual Lasso. In each setting we use 200 samples.
Features Lasso NN NIMO LassoNet Contextual Lasso
Vanilla regr. 10 0.010 0.939 0.048 0.015 -
Toy Example 3 18.982 0.446 0.166 0.628 0.646
Setting 1 5 3.164 1.109 0.030 0.078 0.381
Setting 2 10 3.340 1.482 0.217 2.991 3.418
Setting 3 50 13.122 13.718 0.334 1.342 13.070
Classification Our method can be extended to generalized linear models and here we showcase it for logistic regression. In order to create data for synthetic classification we follow a similar procedure as in the regression case, where we directly control the linear coefficients and the nonlinear interactions. The only difference is that we further apply a link function and binarize the output and get the labels. Note that this way we loose information about the original linear coefficients, which cannot be recovered exactly anymore. The sparsity in the coefficients still remains crucial. Also in this case we show various settings for different dimensionality, nonlinear terms and different number of uninformative features. We provide more details about each setting in the Appendix. In Table 2 we report the classification accuracy on the test set for all methods. Despite our efforts, we were not able to train Contextual lasso on such classification settings, even when all features were given as contextual features. Overall, our model significantly outperforms logistic regression on all settings. When analyzing the learned coefficient, which for brevity we report in the Appendix, we can see that only NIMO recovers the correct sparsity patterns. As argued before, this allows to learn the correct nonlinear interactions. Other nonlinear models and hybrid approaches perform similarly to NIMO but fail to provide the interpretability of the coefficients as in a linear model.
Table 2: Classification accuracy for synthetic classification settings (see the Appendix for more details). We compare with logistic regression, NN and LassoNet. In each setting we use 200 samples.
Features Log. Regr. NN NIMO LassoNet
Setting 1 3 0.59 0.90 0.92 0.89
Setting 2 10 0.68 0.83 0.85 0.82
Setting 3 50 0.74 0.69 0.84 0.90
Figure 5: Diabetes datasets results. Left: MSE loss on test set. All methods achieve comparable results. Middle: Learned linear coefficients by Lasso and NIMO. NIMO selects sparser coefficients (“s1, “s2”, “s4”). Right: the first fc layer of NIMO is very sparse; few nonlinearities are relevant. Errorbars are obtained by running all models across 5 different datasets splits.
4.3 Real Datasets
We further evaluated our model on two real-world datasets: the diabetes dataset (Efron et al., 2004) and the Boston housing dataset (Belsley et al., 2005). In all the experiments, we use a 60%-20%-20% train-validation-test split. We compare our model with Lasso regression, a standard neural network and LassoNet (Lemhadri et al., 2021) in terms of MSE loss and feature sparsity.
@@ 4.3 Real Datasets
Boston dataset
The Boston housing dataset contains 506 instances, each with 13 features, and the target is the median value of owner-occupied homes. As shown in Figure 6, our method performs significantly better than Lasso regression and shows compatible MSE test loss with the other nonlinear and hybrid approaches. This suggests that nonlinear interactions are crucial to explain the data. While both the neural network and LassoNet are able to capture those nonlinearities, they both lack the interpretability provided by our model. It is important to note that the Boston housing dataset has been criticized for containing controversial features  [^1] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html. The feature of interest is termed “B”. From the coefficient plot in Figure 6 (middle), we observe that the significance of the feature “B” is greatly reduced in our model compared to Lasso regression. This surprising result suggests that the reason why such feature appeared to be relevant is due to some artifacts of linear models. Thanks to the decoupling between the linear coefficients and the nonlinearities in NIMO, our results suggests that the role of the feature “B” towards the prediction might be irrelevant. Note that the feature “B” is also not relevant for the nonlinear terms, as shown in Figure 6.
Figure 6: Boston datasets results. Left: Comparison of MSE losses. NIMO performs on par with other methods and significantly outperforms Lasso, suggesting nonlinear interactions are crucial. Middle: Learned linear coefficients by Lasso and NIMO. The two models learn coefficient with a comparable sparsity. Relevantly, the feature “B” is almost irrelevant for NIMO. Right: sparsity in the first fc layer of NIMO. Errorbars are obtained by running all models over 5 datasets splits.

### v2

@@ 3.3 Training via parameter elimination
Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\mathbf{y}\in\mathbb{R}^{n}$. As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$. The model in Eq. 2 can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication. Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$. To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon & Moolgavkar, 1988; Murphy & Van der Vaart, 2000). The key idea is to eliminate $\bm{\beta}$ by first solving for its closed-form expression in terms of $\bm{u}$, and then substituting it back into the objective. This reduces the problem to an optimization over $\bm{u}$ only:
$\min_{\bm{\beta},\bm{u}}\mathcal{L}(\bm{\beta},\bm{u})\;\;\longrightarrow\;\;\min_{\bm{u}}\mathcal{L}(\hat{\bm{\beta}}(\bm{u}),\bm{u}),\quad\text{where}\quad\hat{\bm{\beta}}(\bm{u})=(B_{\bm{u}}^{T}B_{\bm{u}}+\lambda I)^{-1}B_{\bm{u}}^{T}\mathbf{y}.$ (5)
We can then efficiently optimize the objective with gradient descent over $\bm{u}$ only. Furthermore, within this framework we can incorporate sparsity by replacing the $\ell_{2}$ penalty with the $\ell_{1}$ penalty:
@@
Unlike ridge regression, Lasso does not admit a closed-form solution for $\bm{\beta}$. To eliminate $\bm{\beta}$ in this setting, we use adaptive ridge regression, which admits at each step a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and is equivalent to Lasso at the optimum (Grandvalet, 1998); see Appendix A for a detailed proof. This enables us to efficiently optimize the problem using gradient descent while still enforcing sparsity. The full procedure is summarized in Algorithm 1 in the Appendix. A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan & Li, 2001). In some applications, this can lead to significant bias and degrade model performance. To mitigate this issue, we can extend the $\ell_{1}$ norm regularization to a sub-$\ell_{1}$ pseudo-norm, which penalizes large coefficients less aggressively while still promoting sparsity. A detailed introduction to this approach is provided in Appendix B.3.
@@
The same idea can be extended to any generalized linear model of the form: $f(\mathbb{E}[Y|\mathbf{x}])=\beta_{0}+\beta_{1}x_{1}+\dots+\beta_{p}x_{p}=\mathbf{x}^{T}\bm{\beta}$, where $f$ is a link function. For simplicity, consider logistic regression with $\ell_{1}$ penalty. As before, we first get a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$. What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh & Nelder, 1989), which is a weighted least-squares approximation of the original problem:
@@
where $W_{ii}=\sigma(\eta_{i})\bigl[1-\sigma(\eta_{i})\bigr]$ and $\bm{\eta}=B_{\bm{u}}\bm{\beta}$. The so-called working response $\bm{z}$ is computed as $\bm{z}=\bm{\eta}+W^{-1}\bigl(\mathbf{y}-\sigma(\bm{\eta})\bigr)$. Eq. 7 is also a Lasso problem and we can employ adaptive ridge regression to find a closed form expression for $\hat{\bm{\beta}}(\bm{u})$. Then, we substitute $\hat{\bm{\beta}}(\bm{u})$ back and optimize over the neural networks parameters $\bm{u}$ via gradient descent. Appendix B.2 includes a detailed derivation.
3.4 Proof-of-concept: A toy example
We illustrate NIMO on a simple three-dimensional toy example. The data is generated as:
where $\epsilon\sim\mathcal{N}(0,0.1^{2})$ is Gaussian noise. This example involves three features $\{x_{1},x_{2},x_{3}\}$, but $x_{3}$ is inactive since it is multiplied by a zero coefficient. We generate 400 samples and use a 200-100-100 train-validation-test split. As a baseline, we compare against Lasso regression. We further apply group $\ell_{2}$ regularization to the weight matrix of the first fully connected layer in the neural network to enhance interpretability. Further implementation details are given in Appendix C.2. The results are shown in Figure 3. On the left, we see that the estimated coefficients $\bm{\beta}$, which also represent the marginal effects at the mean, learned by NIMO align with the ground truth; in particular, NIMO correctly identifies $\beta_{3}=0$, i.e. that $x_{3}$ is uninformative. In the middle, the nonlinear corrections $g_{j}(\mathbf{x}_{-j})$ recover the true underlying interactions for both $x_{1}$ and $x_{2}$. This highlights an important interpretability aspect: while $\beta_{1}=3$ and $\beta_{2}=-3$ represent the global population-level effects, the neural network outputs vary across individual inputs, providing per-instance corrections. Therefore, NIMO not only preserves the global interpretability of the coefficients but also captures the correct local deviations at the per-instance level. Finally, the right panel shows that neurons connected to $x_{3}$ in the first fully connected layer are sparse, permitting an additional layer of interpretability: $x_{3}$ does not contribute to nonlinear interactions.
@@ 4 Experiments
In this section, we experimentally evaluate the proposed NIMO, focusing on interpretability through marginal effects. We first introduce the baseline methods and the experimental protocol used for comparison. Next, we formulate three hypotheses and present empirical results that examine both the interpretability and predictive performance of NIMO. Additional results are provided in Appendix D.
4.1 Methods and protocol
Methods:
In this paper, we compare our method against Lasso (Tibshirani, 1996), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024). LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective. NAMs learn a linear combination of neural networks that each attend to a single input feature. IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
Table 1: MSE loss for synthetic regression settings.
Method Setting 1 Setting 2 Setting 3
(#features) (5) (10) (50)
Lasso 3.164 3.340 13.122
NN 1.109 1.482 13.718
NAM 3.427 5.126 16.543
IMN 0.137 1.188 6.308
LassoNet 0.078 2.612 1.738
NIMO 0.024 0.197 0.380
Protocol:
We conduct experiments on both synthetic and real datasets to verify the interpretability and performance of our model. For synthetic datasets, we create them by explicitly controlling the linear coefficients and the nonlinearities, similarly to Eq. 8. We explore multiple nonlinearities, different dimensionalities and also use several uninformative features (i.e. with zero coefficient) to test if we can recover sparse signals. The details about the exact data generating process are provided in the Appendix D.2. For real datasets, we choose several popular and well-studied datasets from UCI Machine Learning Repository  [^1] https://archive.ics.uci.edu/datasets.
4.2 Hypotheses and experiment results
Hypothesis 1:
NIMO outperforms other methods on synthetic datasets and remains robust in low-data regime.
We compare NIMO with the above methods on the synthetic datasets and evaluate predictive performance using mean squared error (MSE). The results are presented in Table 1. In each setting, we use 200 samples for training. In this low-data regime, the naive neural network tends to overfit even with a relatively shallow MLP, while the linear model is too simple to capture the complex underlying nonlinearities. NAM, which assigns a separate MLP to each feature component, is even more prone to overfitting than the naive neural network. Overall, the results show that NIMO not only outperforms the other methods by effectively capturing complex nonlinear feature interactions, but also remains robust in low-data settings due to the regularization and optimization strategies employed.
Hypothesis 2:
NIMO accurately estimates the marginal effects at the mean (MEM), preserving the additive contribution of each feature at the population-level.
@@ 4.2 Hypotheses and experiment results
This experiment evaluates whether NIMO can recover population-level interpretability in terms of marginal effects at the mean. For the other methods that rely on automatic differentiation, MEM can certainly be computed, but they do not yield meaningful interpretations. They reduce to arbitrary numerical values that neither explain the prediction for an instance nor capture population-level effects. In contrast, the specific structure of NIMO ensures that the marginal effects at the mean coincide with the linear coefficients, as in the linear models, providing a clear and intelligible interpretation. Therefore, we restrict the comparison to Lasso only, which directly yields globally interpretable coefficients. As a sanity check, we first test NIMO on a Vanilla regression dataset with purely linear coefficients in $\{-5,-4,\ldots,3,4\}$ and no nonlinearities. This verifies whether the neural network component of NIMO interferes with the linear part when no nonlinear effects are present. As shown in Figure 4 (left), NIMO perfectly recovers the ground truth coefficients, confirming that the neural network does not distort the linear coefficients, maintaining the intelligibility of MEM. We then extend the analysis to three additional regression settings (Setting 1, 2, 3) with varying nonlinearities and sparsity levels (see Appendix D.2 for details). Unlike Lasso, NIMO not only recovers the true coefficients but also correctly identifies the sparsity pattern, i.e., which features are uninformative (Figure 4, bottom-right). Based on these results, we conclude that NIMO estimates the accurate marginal effects at the mean and preserves the additive population-level contribution of each feature.
Hypothesis 3:
NIMO achieves a favorable trade-off between interpretability and predictive performance in practice.
Figure 5: Predictive performance (MSE) the diabetes, Boston housing and superconductivity dataset.
This experiment aims to evaluate whether NIMO can achieve strong predictive performance on real datasets and how the results can be interpreted. We select several widely used datasets from the UCI Machine Learning Repository: the diabetes dataset (Efron et al., 2004), the Boston housing dataset (Belsley et al., 2005), and the superconductivity dataset (Hamidieh, 2018). Each dataset is randomly shuffled and split multiple times with different seeds (9 seeds for the diabetes dataset and the Boston housing dataset, 5 seeds for the superconductivity dataset), and the experiments are performed once on each split using identical model settings.
(a) Diabetes
(b) Boston
Figure 6: MEMs for Lasso and NIMO on the diabetes (left) and Boston housing dataset (right).
We first examine the predictive performance of NIMO. The box plot of the mean squared error for each method is shown in Figure 5. NIMO generally achieves performance compatible with the other methods and attains the best performance on the superconductivity dataset. To assess interpretability, we first visualize the marginal effects at the mean (MEM) obtained from Lasso and NIMO for the diabetes and Boston housing datasets in Figure 6. Recall that MEM estimates marginal effects at the average values of the input features, reflecting the additive population-level contribution of each feature. The visualizations reveal that the MEM from NIMO differ substantially from those of Lasso, reflecting NIMO’s ability to capture nonlinear feature interactions that Lasso cannot represent. For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix.
We then use MEM to rank the features on the diabetes and Boston housing datasets. For comparison, we also apply the post-hoc explainer SHAP (Lundberg & Lee, 2017) to our model and rank features according to the global SHAP values. The results are shown in Table 2. On the diabetes dataset, all methods behave similarly in identifying the most important features, which is expected since the data lacks complex nonlinear interactions. This is further supported by the comparable predictive performance across methods, as shown in Figure 5 (left). Notably, SHAP produces identical feature rankings to MEM for NIMO. In contrast, for the Boston housing dataset, the methods yield inconsistent feature rankings, suggesting that nonlinear interactions play a crucial role in explaining the data. Although MEM can also be computed for other nonlinear methods, they do not provide the same level of intelligible interpretability as our model. For NIMO, SHAP rankings are largely consistent with MEM, though not perfectly aligned. The most pronounced difference appears with LassoNet, which assigns high importance to features considered least relevant by other methods, such as INDUS. We attribute this to how LassoNet handles nonlinearities.
Table 2: Feature rankings according to MEMs on the diabetes and Boston housing datasets. A lower ranking is associated with a higher feature importance. Features whose rankings are consistently similar across multiple methods are highlighted in yellow.
@@
Feature SHAP NIMO LassoNet NAM IMN Lasso age 7 7 9 5 9 10 sex 5 5 5 7 5 5 bmi 1 1 2 2 2 2 bp 3 3 3 3 3 4 s1 10 10 7 10 8 3 s2 9 9 10 9 7 6 s3 4 4 4 6 6 8 s4 8 8 8 8 4 7 s5 2 2 1 1 1 1 s6 6 6 6 4 10 9
Boston Housing Dataset
Feature SHAP NIMO LassoNet NAM IMN Lasso CRIM 10 9 3 5 8 9 ZN 13 13 10 11 12 8 INDUS 11 12 2 12 13 13 CHAS 12 8 5 8 10 11 NOX 8 11 9 10 9 7 RM 2 2 8 6 1 3 AGE 6 7 6 9 5 12 DIS 3 3 1 3 3 2 RAD 5 1 12 13 7 4 TAX 4 5 11 7 6 6 PTRATIO 7 6 4 4 4 5 B 9 10 13 2 11 10 LSTAT 1 4 7 1 2 1
4.3 Ablation study
To verify the effectiveness of masking out the $j$-th component of an instance when passing it through the neural network, we conducted an ablation study. Specifically, we examined the effect of masking out the $j$-th component as well as the effect of removing position encoding. From the ablation results across the three settings (Figure 7, Figure 15, and Figure 16), we observe several consistent patterns:
  • Without masking the $j$-th component, the first fc layer of the neural network can still exhibit sparsity patterns, but they are less distinct than when masking is applied. We attribute this to the position encoding, which enables the network to identify the feature index.
  • Without masking the $j$-th component, the MEMs are confounded by nonlinear interactions, so their interpretation as additive regression coefficients and the meaning of their magnitudes are lost.
  • Further removing the position encoding distorts the sparsity patterns in the first fc layer and alters the magnitude of the coefficients, undermining interpretability.
Figure 7: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 1.